Gender Daily Level of Exercise Much Moderate Little Total Male 12 57 56 125 Female 36...

Gender

Daily Level of Exercise

Much

Moderate

Little

Total

Male

12

57

56

125

Female

36

150

81

267

Total

48

207

137

392

1) For five women drawn at random from the student sample, the chance that none of them exercise only a little is approximately:

2) The Relative Risk of little exercise for male students relative to female students is approximately:
Answer #1

a)

hyper geometric distribution

N=667

k=81

n=5

x=0

P(X=0) = [(C(81,0)*C(667-81,5) / C(667,5) ] = 0.52234

b)

relative risk  of little exercise for male students relative to female students = (56/125)/(81/267)=1.4767